if 5 apples in a barrel of 25 apples are rotton, what is the expected number of rotton apples in a sample of 2 apples?
Answer by stanbon(75887)

if 5 apples in a barrel of 25 apples are rotten, what is the expected number of rotten apples in a sample of 2 apples?
----
Binomial Problem with n = 2 and p(rotten) = 5/25 = 1/5
----
Expectec number = mean = np = 2*(1/5) = 2/5
====================
Cheers,
Stan H.
===========